OpenStudy (anonymous):

For the function f(x) = |2x - 7| - 3, find f(2).

OpenStudy (anonymous):

lets say our function is f(x) = x+1 if we're trying to solve for f(1) then f(1) = 1 +1 if we're trying to solve for f(2) then f(2) = 2+1 if we're trying to solve for f(-100) then f(-100) = -100+1 see the pattern?

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

do you think you can do it by yourself now?

OpenStudy (anonymous):

yeah i got 8 is that right?